OpenStudy (anonymous):

Solve: (4x - 5)(3x + 2) =0 {5, -2} {5/4, - 2/3} {-5/4, 2/3} no solution

OpenStudy (hoblos):

(4x - 5)(3x + 2) =0 so 4x-5=0 or 3x+2=0 x=5/4 x=-2/3